At what temperature do Celsius and Fahrenheit are equivalent in value? Prove your answer.

1
Expert's answer
2022-01-05T14:30:19-0500

Celsius and Fahrenheit are equivalent at -40 degrees (Celsius or Fahrenheit). Let's prove that. Let's first use the conversion formula from degrees Fahrenheit to degrees Celsius:


"^{\circ}C=\dfrac{5}{9}(^{\circ}F-32),""x=\dfrac{5}{9}(x-32),""x-\dfrac{5}{9}x=-17.778,""0.444x=-17.778,""x=\dfrac{-17.778}{0.444}=-40\ degrees."

Let's use the conversion formula from degrees Celsius to degrees Fahrenheit:


"^{\circ}F=(^{\circ}C\times\dfrac{9}{5})+32,""x=(x\times\dfrac{9}{5})+32""x-\dfrac{9}{5}x=32,""-0.8x=32,""x=\dfrac{32}{-0.8}=-40\ degrees."

Therefore, we prove that Celsius and Fahrenheit are equivalent at -40 degrees (Celsius or Fahrenheit).